OpenStudy (anonymous):

Help really bad at fractions For her science project, Sylvia collected flowers and measured their length. The line plot displays the lengths of the flowers. If the flowers were laid end to end, what is the total length of all the flowers? Length of Flowers A. 5 1/4 feet B. 5 3/4 feet C. 5 1/2 feet D. 5 5/8 feet

OpenStudy (anonymous):

https://app120.studyisland.com/PICS/158902_15a.png

OpenStudy (anonymous):

\(\dfrac{1}{8} + \dfrac{1}{4} + \dfrac{3}{8} + \dfrac{3}{8} + \dfrac{3}{8} + \dfrac{3}{8} + \dfrac{1}{2} + \dfrac{1}{2} + \dfrac{5}{8} + \dfrac{5}{8} + \dfrac{3}{4} + \dfrac{7}{8}\) We can change all the fractions to denominators of 8. \(\dfrac{1}{4} \rightarrow \dfrac{1 \times 2}{4 \times 2} \rightarrow \dfrac{2}{8}\) \(\dfrac{1}{2} \rightarrow \dfrac{1 \times 4}{2 \times 4} \rightarrow \dfrac{4}{8}\) \(\dfrac{3}{4} \rightarrow \dfrac{3 \times 2}{4 \times 2} \rightarrow \dfrac{6}{8}\) Plug these back in: \(\dfrac{1}{8} + \dfrac{2}{8} + \dfrac{3}{8} + \dfrac{3}{8} + \dfrac{3}{8} + \dfrac{3}{8} + \dfrac{4}{8} + \dfrac{4}{8} + \dfrac{5}{8} + \dfrac{5}{8} + \dfrac{6}{8} + \dfrac{7}{8}\) Now we can just add them all up. \(\dfrac{1 + 2 + 3 + 3 + 3 + 3 + 4 + 4 + 5 + 5 + 6 + 7}{8}\) So can you add 1 + 2 + 3 + 3 + 3 + 3 + 4 + 4 + 5 + 5 + 6 + 7?

OpenStudy (anonymous):

\[46/8\]

OpenStudy (anonymous):

but i need that in feet

OpenStudy (anonymous):

I know, so now we divide: